OpenStudy (anonymous):

simplify. [9^4*x^(3/2)*y(5/2)]/[81^6*x^-.5*y^-1] show steps

OpenStudy (anonymous):

\[9^4x^{3/2}y^{5/2}/81^6x^{-1/2}y^{-1}\]

OpenStudy (anonymous):

Rules a^b * a^c = a^{b+c} a^b/a^c =a^{b-c} (a^b)^c = a^{bc} 9^4 is 81^2 and follow these rules (the second one) to get your answer